OpenStudy (anonymous):

This one is different from the last. Call for solution for x & y. (2x+3i)-(6+2yi)=5-3i

OpenStudy (anonymous):

(2x+3i)-(6+2yi)=5-3i=(2x-6) + (3+2y)i=5-3i..so 2x-6=5 and 3+2y=-3 ..solving x=11/2 and y=-3

OpenStudy (anonymous):

Where did the i go?

OpenStudy (anonymous):

equating real and complex parts....i indicates complex number